OpenStudy (anonymous):

A piece of wire 5 inches long is to be cut into two pieces. One piece is x inches long and is to be bent into the shape of a square. The other piece is to be bent into the shape of a circle. Find an expression for the total area made up by the spare and the circle as a function of x.

OpenStudy (anonymous):

Area of a square is \[s ^{2}\] The bent piece of wire will have side lengths of x/4. So the area of the square is \[(x/4)^{2}\] The rest of the wire is 5" - x. That will comprise the circumference of the circle. \[C = 2 \pi r\] We can solve this for r and then use that to find the are of the circle. \[r = (5-x)/2 \pi\] Area of a circle is \[\pi r ^{2}\] So you can substitute the r value into the equation and solve for that area as well. Enjoy.